prestonmckenzie:

The height and the base of a parallelogram are in the ratio 5 : 6. If the area of the parallelogram is 1080 sq. cm. Determine its height and base.

YourDaddy6969:

wHaT

Sock:

@yourdaddy6969 wrote:
wHaT
please do not respond to a question without the intention of helping/explaining the answer to the person. There was a whole post made about this: https://questioncove.com/study#/updates/616b3dba114cc78f1acf9700

surjithayer:

let the base=6x height=5x area=base *height \[6x \times 5x=1080\] \[x^2=\frac{ 1080 }{ 30 }=36\] \[x=\sqrt{36}=6 (taking~only~positive~value.)\] so base=6x=6*6=36 cm height=6*5=30 cm

YourDaddy6969:

@sock wrote:
@yourdaddy6969 wrote:
wHaT
please do not respond to a question without the intention of helping/explaining the answer to the person. There was a whole post made about this: https://questioncove.com/study#/updates/616b3dba114cc78f1acf9700
OKay it was a joke but sure

prestonmckenzie:

@surjithayer thank you